# Compression with Global Guidance: Towards Training-free High-Resolution MLLMs Acceleration

### v2

@@ I Introduction
To address the challenges, we specifically design an effective token compression method for high-resolution MLLMs, termed as GlobalCom2 (short for “Global Compression Commander”). The motivation behind our method is that visual tokens derived from the global thumbnail and those from local crops serve distinct functions. While tokens from the thumbnail provide overarching information, the tokens from local crops offer more nuanced details. Therefore, our GlobalCom2 adheres to a “global-to-local” guiding principle as the design philosophy. Specifically, the compression process can be divided into two stages. In the first stage, due to the varying amounts of semantic information contained in local crops, GlobalCom2 uses the information provided by the thumbnail to assess the importance of each crop and thereby allocate different token retention rates to each. In the second stage, GlobalCom2 performs token compression on both the thumbnail and the crops, where the compression within the crops is guided by the thumbnail information to maximize the preservation of visual detail information. In this way, GlobalCom2 adaptively preserves more essential and informative local details during token compression.
@@ III Methodology
In this section, we present our GlobalCom2 in detail. First, we briefly revisit the LLaVA-NeXT model as the most representative high-resolution MLLM in Section III-A. Then, we elaborate on the design of our training-free token compression method GlobalCom2 in Section III-B. Finally, we provide an in-depth analysis of theoretical complexity in Section III-C.
@@ III-A Preliminary: LLaVA-NeXT
Local crops offer fine-grained visual insights specific to different areas of an image, while the global thumbnail acts as a comprehensive visual extractor, capturing and conveying the essential features of each sub-image along with the overall image context.
@@ III-B Global Compression Commander: GlobalCom2
Based on the above analysis, we aim to enhance the inference efficiency of high-resolution MLLMs by directly reducing the computational costs through visual token sequence compression (i.e., token compression). Given that LLaVA-NeXT partitions input images into a global thumbnail followed by a series of local crops as shown in the top-left of Figure 2, performing visual token compression for LLaVA-NeXT involves compressing tokens from both the global thumbnail and local crops. As analyzed in Section III-A, the global thumbnail and local crops serve distinct roles in visual understanding within LLaVA-NeXT, suggesting that different token compression strategies should be applied to each component. Furthermore, since the global thumbnail acts as an “abstractor” of the entire input image’s visual information with a holistic perspective, we propose to leverage it as a “Global Compression Commander” (GlobalCom2) to guide token compression for high-resolution MLLMs, as shown in Figure 2. In what follows, we will elaborate on how our GlobalCom2 guides token compression for both the global thumbnail and local crops.
@@
Since the AnyRes strategy partitions high-resolution images into multiple local crops, the visual token sequences of local crops are significantly longer than the global thumbnail and contain more visual redundancy. Furthermore, each local crop contains distinct visual information, leading to varying amounts of informational content across crops. As shown in Figure 2, the upper two crops contain significant entity information, such as football players, while the lower two crops exhibit considerable visual redundancy, predominantly showing large areas of grass. The visualization in the upper middle of Figure 2 displays the attention values between tokens in the global thumbnail and the [CLS] token, clearly demonstrates higher attention values in the upper regions, indicating that the upper two local crops contain richer semantic information from a global perspective.
Based on the above analysis, we argue that each local crop contributes differently to the overall visual understanding of high-resolution MLLMs, thus warranting varying degrees of token compression for different local crops. Local crops containing rich semantic information in the context of the entire image should retain more visual tokens, providing LLMs with semantically dense visual information to capture crucial local visual details. Conversely, crops with limited semantic content should undergo more aggressive token compression, allowing LLMs to focus on semantically rich visual signals. Therefore, we aim to guide the token compression of local crops from a global perspective. Specifically, GlobalCom2 incorporates two aspects of global guidance for local crop token compression: retention ratio allocation and token importance evaluation.
@@ IV Experiments
TABLE I: Comparison with other training-free MLLM token compression methods with LLaVA-NeXT-7B across 10 benchmarks. “Average” represents the mean performance rate between token compression methods and vanilla LLaVA-NeXT-7B across all 10 benchmarks. We evaluate the methods at decreasing retention ratios from 75% to 10%, with the best results highlighted.
@@
TABLE II: Comparison with other training-free MLLM token compression methods with LLaVA-NeXT-13B across 10 benchmarks.
@@
IV-A Experimental Setting
Benchmark Details
Implementation Details
Comparison Details
IV-B Main Results
@@
Table I presents the performance comparison of FastV [20], SparseVLM [19], FasterVLM [21], and our GlobalCom2 across 10 benchmarks using LLaVA-NeXT-7B under various retention ratios. The results reveal several key findings:
@@
Given that FasterVLM outperforms FastV and SparseVLM in Table I significantly, and due to the unavailability of FastV and SparseVLM results on LLaVA-NeXT-13B, Table II focuses on comparing our GlobalCom2 with FasterVLM. The comparison results in Table II highlights several observations:
@@ IV-C Ablation Study and Analysis
TABLE III: Ablation of significance criteria of local crops.
@@
Table III shows the impact of different retention ratio allocation strategies for local crops. We adopt four different strategies:
@@
As shown in Table III, our adopted Softmax (sum) strategy achieves the best performance, maintaining 96.1% performance of the vanilla LLaVA-NeXT-7B. Despite "Uniform", the other strategies acknowledge varying importance across local crops. However, both $\mathbf{n_{\text{top-{k}}}}$ and Softmax(max) strategies only consider the strongest visual representation capability within each local crop, rather than evaluating the overall importance of the entire local crop to the global image. In contrast, our adopted Softmax (sum) strategy adaptively allocates retention ratios based on each local crop’s importance to the global image, thereby preserving more semantically visual information from high-resolution images and enabling LLM to better capture fine-grained visual signals.
TABLE IV: Ablation of token retention criteria of local crops.
@@
Table IV further examines the impact of different local crop token retention strategies on LLaVA-NeXT token compression, where “global” and “local” refer to token importance consideration at global and local levels, respectively. We observe that considering only global importance yields lower performance than considering only local importance. This is because the global-only approach overlooks important fine-grained visual details in local regions, leading to degraded performance on tasks that rely heavily on visual details, such as the text recognition benchmark VQAT. Our GlobalCom2 achieves optimal performance by jointly considering token importance at both global image and local crop levels.
@@
TABLE V: Efficiency analysis of our GlobalCom2 with LLaVA-NeXT-7B on one NVIDIA A100-SXM4-80GB GPU.
Upper Bound, 2880 Tokens
LLaVA-NeXT-7B 41.65 107 1536 170.7 15:35 1519.0
Ratio=75%, Retain up to 2160 Tokens
GlobalCom2 30.42 (↓26.96%) 70.7 (↓33.93%) 1126.4 (↓26.65%) 119.9 (↓29.79%) 12:17 (↓21%) 1548.4
Ratio=50%, Retain up to 1440 Tokens
GlobalCom2 19.74 (↓52.61%) 43.1 (↓59.72%) 755 (↓50.85%) 74.5 (↓56.38%) 9:21 (↓40%) 1552.9
Ratio=25%, Retain up to 720 Tokens
GlobalCom2 9.60 (↓76.95%) 23.9 (↓77.66%) 377 (↓75.45%) 34.6 (↓79.73%) 7:27 (↓52%) 1493.5
Ratio=10%, Retain up to 288 Tokens
GlobalCom2 3.77 (↓90.95%) 16.5 (↓84.58%) 151 (↓90.17%) 13.3 (↓92.21%) 5:48 (↓63%) 1365.5
TABLE VI: Efficiency analysis of our GlobalCom2 with LLaVA-NeXT-13B on one NVIDIA A100-SXM4-80GB GPU.
Upper Bound, 2880 Tokens
LLaVA-NeXT-13B 80.0 172 2457.6 295.1 20:39 1575.9
Ratio=75%, Retain up to 2160 Tokens
GlobalCom2 58.7 (↓26.62%) 116 (↓32.56%) 1843.2 (↓25.03%) 211.6 (↓28.29%) 16:30 (↓20%) 1567.9
Ratio=50%, Retain up to 1440 Tokens
GlobalCom2 38.3 (↓52.12%) 72.4 (↓57.91%) 1228.8 (↓50.00%) 134.6 (↓54.39%) 12:49 (↓38%) 1553.5
Ratio=25%, Retain up to 720 Tokens
GlobalCom2 18.7 (↓76.62%) 42.4 (↓75.35%) 590 (↓76.00%) 64.1 (↓78.27%) 8:48 (↓57%) 1531.2
Ratio=10%, Retain up to 288 Tokens
GlobalCom2 7.4 (↓90.75%) 30.8 (↓82.09%) 236 (↓90.40%) 25.0 (↓91.53%) 6:53 (↓67%) 1399.5
@@
We conduct a comprehensive analysis of both theoretical and practical efficiency of our GlobalCom2 with LLaVA-NeXT-7B/13B on a single NVIDIA A100-SXM4-80GB GPU, as shown in Table V and Table VI. For theoretical efficiency, metrics except "TFLOPs" and "Total Time" are estimated using LLM-Viewer [45]. Given that the sequence length of visual tokens substantially exceeds that of textual and system tokens, we exclude the latter two from our analysis. For practical efficiency, "Total Time" and the corresponding "Performance" measurements are conducted on the MME benchmark [41].
As demonstrated in Table V and Table VI, our GlobalCom2 significantly enhances the computational efficiency of LLaVA-NeXT models in both theoretical and practical aspects. While maintaining comparable model performance, it substantially reduces GPU memory utilization and considerably accelerates inference speed. Specifically, our GlobalCom2 achieves remarkable efficiency improvements: reducing computational overhead by nearly 91%, GPU memory usage by approximately 90%, KV-Cache storage by around 90%, prefill time by around 92%, and inference time by around 65%, while maintaining over 90% of the original performance by LLaVA-NeXT. Notably, these efficiency gains are accomplished without any additional re-training, highlighting the practical utility of our GlobalCom2.
@@ IV-D Visualizations
To evaluate the effectiveness of global guidance in local crop token compression, we present token compression visualizations on the VQAT benchmark, which effectively measures the performance of token compression in high-resolution MLLMs, as shown in Figure 4.
It is clear that our GlobalCom2 effectively removes redundant regions in both global thumbnails and local crops, demonstrating its ability to preserve entity-rich image regions and provide semantically meaningful visual signals to LLMs. When certain local crops exhibit significant visual redundancy within a complete image (as shown in the top-right case where the last crop largely consists of a camera’s golden casing), GlobalCom2 effectively assigns lower retention ratios to these semantically sparse crops to eliminate redundant regions. Conversely, for crops with rich semantic content (such as the text-heavy first crop in the top-right case), GlobalCom2 allocates higher retention ratios to preserve these crucial visual tokens for high-resolution image understanding.

### v3

@@ I Introduction
In this paper, we perform systematic experiments and analysis to examine the roles of global thumbnails and local crops in high-resolution MLLM image understanding capabilities. Through our empirical investigation, we identify two key insights: (1) Visual tokens from global thumbnails and local crops serve complementary purposes, where thumbnail tokens capture holistic context while local crop tokens preserve fine-grained visual details. (2) Visual tokens from different local crops exhibit varying levels of semantic richness and redundancy, which directly impacts their contributions to high-resolution MLLM image understanding.
Building on these insights, we propose an effective token compression method GlobalCom2 (short for “Global Compression Commander”), specifically designed for high-resolution MLLMs. Our method follows a “global-to-local” design philosophy and operates in two stages. In the first stage, GlobalCom2 leverages thumbnail information to evaluate the semantic significance of each local crop, enabling adaptive token retention rates across different regions. In the second stage, it performs token compression on both the thumbnail and local crops, where the compression of local crops is guided by thumbnail information to maximize the preservation of crucial visual details. Through this hierarchical approach, GlobalCom2 achieves efficient and adaptive token compression while maintaining essential visual information.
@@ III Methodology
In this section, we present our GlobalCom2 in detail. First, we briefly revisit LLaVA-NeXT, a representative high-resolution MLLM, in Section III-A. Through systematic analytical experiments, we identify fundamental insights into the characteristics of high-resolution MLLMs and, based on these insights, present the detailed design of our training-free method GlobalCom2 in Section III-B. Finally, we provide an in-depth analysis of theoretical complexity in Section III-C.
@@ III-B Global Compression Commander: GlobalCom2
TABLE I: Performance analysis of LLaVA-NeXT under different image inputs across various benchmarks.
Method VQAT POPE SQA MMB MM-Vet
LLaVA-NeXT-7B 61.3 86.5 70.1 67.4 43.9
(a) Effects of Global Thumbnail and Local Crops
Global thumbnail only 56.8 84.6 70.9 67.3 42.4
Local crops only 57.5 87.3 68.0 65.0 40.8
(b) Effects of Different Local Crops
Random crop removal 56.4 86.1 66.9 64.7 36.6
Important crop removal 54.3 85.3 65.3 65.4 36.1
Unimportant crop removal 57.5 87.0 68.4 66.3 38.5
Based on the above analysis, we aim to enhance the inference efficiency of high-resolution MLLMs by directly reducing the computational costs through visual token sequence compression (i.e., token compression). Given that LLaVA-NeXT partitions input images into a global thumbnail followed by a series of local crops as shown in the top-left of Figure 2, performing visual token compression for LLaVA-NeXT involves compressing tokens from both the global thumbnail and local crops.
We begin with an exploratory experiment to investigate how global thumbnails and local crops differently contribute to image understanding in high-resolution MLLMs by using them separately as input to LLaVA-NeXT. As shown in Table I (a), using global thumbnail alone yields superior performance on general visual perception benchmarks such as SQA [35], MMB [36], and MM-Vet [37]. This advantage stems from global thumbnails serving as comprehensive visual extractor, providing holistic visual information through complete ViT encoding. In contrast, using only local crops, where ViT independently encodes each crop, shows inferior performance on these general visual perception tasks due to the lack of holistic visual information. However, local crops excel in fine-grained perception tasks like VQAT (requiring text recognition in images) and POPE (testing model hallucination), as they provide more detailed visual features. Therefore, we argue that global thumbnails and local crops play distinct roles in visual perception tasks of high-resolution MLLMs, thus requiring different token compression strategies.
Furthermore, given the global thumbnail’s capability as a comprehensive visual extractor that captures the entire input image’s visual information from a holistic perspective, we propose to leverage it as a “Global Compression Commander” (GlobalCom2) to guide token compression for high-resolution MLLMs, as shown in Figure 2. In what follows, we will elaborate on how our GlobalCom2 guides token compression for both the global thumbnail and local crops.
@@
Due to the AnyRes strategy’s image partitioning, local crops produce longer visual token sequences with more redundancy compared to the global thumbnail. Furthermore, each local crop contains distinct visual information, leading to varying amounts of informational content across crops. As shown in Figure 2, the upper two crops contain significant entity information, such as football players, while the lower two crops exhibit considerable visual redundancy, predominantly showing large areas of grass. The visualization in the upper middle of Figure 2 displays the attention values between tokens in the global thumbnail and the [CLS] token, clearly demonstrates higher attention values in the upper regions, indicating that the upper two local crops contain richer semantic information from a global perspective.
Moreover, based on our previous analysis, the attention scores with the [CLS] token effectively measure the semantic information contained in a token. Therefore, we partition the 2D attention map of the global thumbnail according to the crop segmentation pattern, and sum up the scores in each sub-attention map to obtain the global importance of that crop. We investigate the varying contributions of different local crops to image understanding in high-resolution MLLMs by selectively dropping local crops from the input image sequence of LLaVA-NeXT (randomly dropping one crop, dropping the most important crop, and dropping the least important crop, respectively). As demonstrated in Table I (b), different local crops contribute differently to visual understanding due to their varying semantic richness. Specifically, when dropping the most important versus the least important crops, we observe significant performance gaps across various visual understanding tasks, particularly in VQAT for text recognition, where the gap reaches 3.2. This indicates that different crops contribute varying degrees of detail comprehension, while also suggesting that semantically richer local crops are more crucial for fine-grained visual understanding.
Based on the above qualitative and quantitative analysis, we argue that each local crop contributes differently to the overall visual understanding of high-resolution MLLMs, thus warranting varying degrees of token compression.
Local crops containing rich semantic information in the context of the entire image should retain more visual tokens, providing LLMs with semantically dense visual information to capture crucial local visual details. Conversely, crops with limited semantic content should undergo more aggressive token compression, allowing LLMs to focus on semantically rich visual signals. Based on the above analysis and motivated by observations -, we aim to guide the token compression of local crops from a global perspective. Specifically, GlobalCom2 incorporates two aspects of global guidance for local crop token compression: retention ratio allocation and token importance evaluation.
@@ IV Experiments
TABLE II: Comparison with other training-free MLLM token compression methods with LLaVA-NeXT-7B across 10 benchmarks. “Average” represents the mean performance rate between token compression methods and vanilla LLaVA-NeXT-7B across all 10 benchmarks. We evaluate the methods at decreasing retention ratios from 75% to 10%, with the best results highlighted.
@@
IV-A Experimental Setting
Benchmark Details
Implementation Details
Comparison Details
IV-B Main Results
TABLE III: Comparison with other training-free MLLM token compression methods with LLaVA-NeXT-13B across 10 benchmarks.
@@ IV-B Main Results
Table II presents the performance comparison of FastV [20], SparseVLM [19], FasterVLM [21], and our GlobalCom2 across 10 benchmarks using LLaVA-NeXT-7B under various retention ratios. The results reveal several key findings:
@@
Given that FasterVLM outperforms FastV and SparseVLM in Table II significantly, and due to the unavailability of FastV and SparseVLM results on LLaVA-NeXT-13B, Table III focuses on comparing our GlobalCom2 with FasterVLM. The comparison results in Table III highlights several observations:
@@ IV-C Ablation Study and Analysis
TABLE IV: Ablation of significance criteria of local crops.
@@
Table IV shows the impact of different retention ratio allocation strategies for local crops. We adopt four different strategies:
@@
As shown in Table IV, our adopted Softmax (sum) strategy achieves the best performance, maintaining 96.1% performance of the vanilla LLaVA-NeXT-7B. Despite "Uniform", the other strategies acknowledge varying importance across local crops. However, both $\mathbf{n_{\text{top-{k}}}}$ and Softmax(max) strategies only consider the strongest visual representation capability within each local crop, rather than evaluating the overall importance of the entire local crop to the global image. In contrast, our adopted Softmax (sum) strategy adaptively allocates retention ratios based on each local crop’s importance to the global image, thereby preserving more semantically visual information from high-resolution images and enabling LLM to better capture fine-grained visual signals.
TABLE V: Ablation of token retention criteria of local crops.
@@
TABLE VI: Efficiency analysis of our GlobalCom2 with LLaVA-NeXT-7B on one NVIDIA A100-SXM4-80GB GPU.
Upper Bound, 2880 Tokens
LLaVA-NeXT-7B 41.7 107.0 1536.0 170.7 15:35 1519.0
Ratio=75%, Retain up to 2160 Tokens
GlobalCom2 30.4 (↓27%) 70.7 (↓34%) 1126.4 (↓27%) 119.9 (↓30%) 12:17 (↓21%) 1548.4
Ratio=50%, Retain up to 1440 Tokens
GlobalCom2 19.7 (↓53%) 43.1 (↓60%) 755.0 (↓51%) 74.5 (↓57%) 9:21 (↓40%) 1552.9
Ratio=25%, Retain up to 720 Tokens
GlobalCom2 9.6 (↓77%) 23.9 (↓78%) 377.0 (↓76%) 34.6 (↓80%) 7:27 (↓52%) 1493.5
Ratio=10%, Retain up to 288 Tokens
GlobalCom2 3.8 (↓91%) 16.5 (↓85%) 151.0 (↓90%) 13.3 (↓92%) 5:48 (↓63%) 1365.5
TABLE VII: Efficiency analysis of our GlobalCom2 with LLaVA-NeXT-13B on one NVIDIA A100-SXM4-80GB GPU.
Upper Bound, 2880 Tokens
LLaVA-NeXT-13B 80.0 172.0 2457.6 295.1 20:39 1575.9
Ratio=75%, Retain up to 2160 Tokens
GlobalCom2 58.7 (↓27%) 116.0 (↓33%) 1843.2 (↓25%) 211.6 (↓28%) 16:30 (↓20%) 1567.9
Ratio=50%, Retain up to 1440 Tokens
GlobalCom2 38.3 (↓52%) 72.4 (↓58%) 1228.8 (↓50%) 134.6 (↓54%) 12:49 (↓38%) 1553.5
Ratio=25%, Retain up to 720 Tokens
GlobalCom2 18.7 (↓77%) 42.4 (↓75%) 590.0 (↓76%) 64.1 (↓78%) 8:48 (↓57%) 1531.2
Ratio=10%, Retain up to 288 Tokens
GlobalCom2 7.4 (↓91%) 30.8 (↓82%) 236.0 (↓90%) 25.0 (↓92%) 6:53 (↓67%) 1399.5
@@
Table V further examines the impact of different local crop token retention strategies on LLaVA-NeXT token compression, where “Global” and “Local” refer to token importance consideration at global and local levels, respectively. We observe that considering only global importance yields lower performance than considering only local importance. This is because the global-only approach overlooks important fine-grained visual details in local regions, leading to degraded performance on tasks that rely heavily on visual details, such as the text recognition benchmark VQAT. Our GlobalCom2 achieves optimal performance by jointly considering token importance at both global image and local crop levels.
@@
We conduct a comprehensive analysis of both theoretical and practical efficiency of our GlobalCom2 with LLaVA-NeXT-7B/13B on a single NVIDIA A100-SXM4-80GB GPU, as shown in Table VI and Table VII. For theoretical efficiency, metrics except "TFLOPs" and "Total Time" are estimated using LLM-Viewer [45]. Given that the sequence length of visual tokens substantially exceeds that of textual and system tokens, we exclude the latter two from our analysis. For practical efficiency, "Total Time" and the corresponding "Performance" measurements are conducted on the MME benchmark [43].
As demonstrated in Table VI and Table VII, our GlobalCom2 significantly enhances the computational efficiency of LLaVA-NeXT models in both theoretical and practical aspects. While maintaining comparable model performance, it substantially reduces GPU memory utilization and considerably accelerates inference speed. Specifically, our GlobalCom2 achieves remarkable efficiency improvements: reducing computational overhead by nearly 91%, GPU memory usage by approximately 84%, KV-Cache storage by around 90%, prefill time by around 92%, and inference time by around 65%, while maintaining over 90% of the original performance by LLaVA-NeXT. Notably, these efficiency gains are accomplished without any additional re-training, highlighting the practical utility of our GlobalCom2.
@@ IV-D Visualizations
To evaluate the effectiveness of global guidance in GlobalCom2 for both global thumbnail and local crop token compression, we present token compression visualizations with a retention ratio of $R=25\%$ on the VQAT benchmark, which effectively measures the compression performance in high-resolution MLLMs, as shown in Figure 4.
It is clear that our GlobalCom2 effectively removes redundant regions in both global thumbnails and local crops, demonstrating its ability to preserve entity-rich image regions and provide semantically meaningful visual signals to LLMs. When certain local crops exhibit significant visual redundancy within a complete image (as shown in the top-right case where the last crop largely consists of a camera’s golden casing), GlobalCom2 effectively assigns lower retention ratios (i.e., $r_{4}=19\%$) to these semantically sparse crops to eliminate redundant regions. Conversely, for crops with rich semantic content (such as the text-heavy first crop in the top-right case), GlobalCom2 allocates higher retention ratios (i.e., $r_{1}=43\%$) to preserve these crucial visual tokens for detailed understanding.